OpenStudy (anonymous):

SAT subject test problem: What is the reciprocal of 8-3i?

OpenStudy (anonymous):

1/(8-3i)

OpenStudy (anonymous):

that is not one of the answer choices. here are the choices: a) 1/8 b)8+3i/55 c)8-3i/55 d)8+3i/73 e)1/11

OpenStudy (anonymous):

@tomo sorry, could you see if you could do this one also?

OpenStudy (anonymous):

reciprocal of i is -i

OpenStudy (anonymous):

so you could simplify my previous expression knowing that

sam (.sam.):

Multiply by the conjugate of it \[\frac{1}{8-3i} \times \frac{8+3i}{8+3i} \\ \\ \frac{8+3i}{64-9i^2} \] We know that \[i^2=-1\] Can you solve it now?

OpenStudy (anonymous):

so.. the answer would be ... 8+3i/73?

sam (.sam.):

yes

OpenStudy (anonymous):

Thank you :)

sam (.sam.):

welcome :)